OpenStudy (anonymous):

CAN SOMEONE PLLLZZZ HELP!!!!!!!!! which factors can u cancle from the numerator and denominator to simplify 12-6x/5x-10

OpenStudy (anonymous):

6(2-x)/5(x-2)=(6/5)*(2-x)/(x-2)=-6/5

OpenStudy (anonymous):

-6/5 isnt an option 6/5 is and 5/6 is also

OpenStudy (anonymous):

A 6/5 B 5/6 C x+2 D x-2

OpenStudy (anonymous):

answer is (x-2) because -6(x-2)/5(x-2)=-6/5 so the factor which simplifies the numerator and denominator is (x-2)

OpenStudy (anonymous):

ok thnx can you help me out with another question